OpenStudy (anonymous):

Given the system of equations, what is the solution? x + 2y = 7 x - 2y = -1 {(-8, -12)} {(3, 2)} {(-4, 6)}

OpenStudy (mathhelp346):

The method I'll be using is substitution. x-2y=-1 +2y +2y x=2y-1 (2y-1)+2y=7 4y-1=7 +1+1 4y=8 y=2 x-2y=-1 x-2(2)=-1 x-4=-1 +4 +4 x=3

OpenStudy (mathhelp346):

So it would be {(3, 2)}